# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

### v2

@@ 3.1 The KAGIN Layer
To achieve its maximal expressivity, the GIN model relies on the MLP architecture and its universal approximator property. Since KAN is also a universal function approximator, we could achieve the same expressive power using KANs in lieu of MLPs. We thus propose the KAGIN model which is defined as follows:
@@
With theoretically-sound KANs (i. e., with arbitrarily complex components), this architecture is exactly as expressive as the vanilla GIN model with arbitrary layer width. While this is not guaranteed with the spline-based implementation with limited grid size, the empirical results in the original paper demonstrate the great expressive power of KANs [23], especially of small models and setting where regularity is expected.
@@ 3.2 The KAGCN Layer
GCN-based architectures have achieved great success in node classification tasks. While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some of the datasets does not facilitate a fair estimation of KANs’ potential in this context. To this end, we also propose a variant of the GCN model. Specifically, we substitute the parameters and ReLU function of the standard GCN [20] model with a single KAN layer (defined in Equation (6)) to obtain the KAGCN layer:
@@
where $\Phi^{(\ell)}$ denotes a single KAN layer. In the familiar matrix formulation, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with self-loops and $\tilde{\mathbf{D}}$ the diagonal degree matrix of $\tilde{\mathbf{A}}$, the node update rule of KAGCN can be written as:
$\mathbf{H}^{(\ell)}=\Phi^{(\ell)}\Big{(}\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(\ell-1)}\Big{)}$ (9)
where the different rows of $\mathbf{H}^{(\ell)}$ store the representations of the different nodes of the graph.
@@ 4 Empirical Evaluation
In this section, we compare the KAGIN and KAGCN models with the GIN and GCN models in the following tasks: node classification, graph classification and graph regression. The code for reproducing the results is available at https://github.com/RomanBresson/KAGNN. All models are implemented with PyTorch [30]. For KAN layers, we rely on a publicly available implementation [^1] https://github.com/Blealtan/efficient-kan.
@@ 4.1 Node classification
Datasets.
To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer [20] and Ogbn-arxiv [16]) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks. The homophilic networks are already split into training, validation and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross validation indices.
@@
For every dataset and model, we tune the values of the hyperparameters. Specifically, we choose the values that perform best on the validation set (i. e., lowest validation error). To find these values, we use the Optuna package [1]. We set the number of iterations of Optuna equal to $100$ trials.
For all models, the learning rate is chosen from $[10^{-3},10^{-2}]$, the number of message passing layers from $\{1,2,3,4\}$, the hidden dimension size from $\{8,9,\ldots,128\}$ and the weight decay from $\{0,0.0005\}$. For KAGIN and KAGCN, we also choose the grid size from $\{3,4,5\}$ and the spline order from $\{1,2,3,4\}$. Once the best hyperparameter values are found, we evaluate the models on the test set. For the homophilic networks, we initialize and train $10$ different models (we use $10$ different random seeds). We then evaluate the $10$ models on the test set and report the average accuracy. For the heterophilic datasets, we tune each model’s hyperparameters within each fold and we report the average accuracy across the $10$ folds.
@@
Dataset KAGIN GIN KAGCN GCN
Cora 0.7620 $\pm$ 0.0077 0.5528 $\pm$ 0.1112
0.7826
$\pm$ 0.0177
0.7566 $\pm$ 0.0718
Citeseer 0.6837 $\pm$ 0.0117 0.4748 $\pm$ 0.0574 0.6409 $\pm$ 0.0185
0.6891
$\pm$ 0.0074
Ogbn-arxiv 0.2450 $\pm$ 0.0806 0.2110 $\pm$ 0.0628
0.2997
$\pm$ 0.0274
0.2777$\pm$ 0.0699
Cornell
0.5243
$\pm$ 0.0813
$0$$\pm$ 0.1414 0.4892 $\pm$ 0.0547 0.4162 $\pm$ 0.0737
Texas 0.5973 $\pm$ 0.0505
0.6054
$\pm$ 0.0516
0.5811 $\pm$ 0.0440 0.5892 $\pm$ 0.0415
Wisconsin
0.5510
$\pm$ 0.0720
0.4961 $\pm$ 0.1532 0.5196 $\pm$ 0.1020 0.5216 $\pm$ 0.0729
Actor
0.2890
$\pm$ 0.0139
0.2817 $\pm$ 0.0110 0.2755 $\pm$ 0.0134 0.2852 $\pm$ 0.0129
@@ Results.
The results are given in Table 1. We can see that KAGIN outperforms GIN on all but one datasets. On some datasets, the difference in performance between the two models is significant. For example, on Citeseer and Cora, KAGIN offers a respective absolute improvement of $20.89\%$ and $20.9\%$ in accuracy over GIN. On the other hand, KAGCN outperforms GCN on $3/7$ considered datasets, albeit being the best in $2/3$ homophilic networks and exhibiting very similar accuracy in $3/4$ heterophilic ones. The KAGIN and GIN models exhibit the best performance in heterophilic networks while the KAGCN and GCN models achieve the highest accuracy on the homophilic. The use of KANs in the GIN architecture brings performance improvements on $6/7$ datasets, but $5/7$ experiments overlap in the confidence intervals, an overlap that is more prevalent with GCN-based models in $6/7$. Overall we can contend that KAN has a more positive impact on the GIN architecture than on the GCN architecture. Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e.g. GCN accuracy with heterophilic networks [41], however, it improves the models in the majority of the results i.e. $5/7$.
With regards to the optimal hyperparameters, we found that KAGIN required a larger grid size on average compared to KAGCN. In general, a grid size of $4$ was chosen in $8/14$ experiments and the predominant spline order was $1$, appearing in $8/14$ experiments while $3$ appears $2$ times and $4$ only once. The number of hidden layers and their sizes varied significantly through datasets but overall KAGIN had a substantially larger average hidden layer size i. e., $75.4$ compared to $48$ for KAGIN, as $128$ was chosen for $3$ datasets compared to only $1$ for KAGCN. This is in contrast to previous findings contending that GIN-based models requires more complex learning procedures, a pattern that seems to withstand with KANs.
Training times.
We present in Table 2 the training time per epoch for different configurations of the KAGIN and GIN models. We observe that for a given number of message passing layers and hidden dimension size, the KAGIN model is computationally more expensive than GIN. If the grid size and spline order hyperparameters of KAGIN are set to $1$, the difference in running time between the two models is very small (less than $0.02$ seconds per epoch for all configurations). However, the complexity of KAGIN increases are the grid size and spline order increase. Overall, our results suggest that the running time of KAGIN is slightly greater than that of GIN, and by no means prohibitive.
Table 2: Training time per epoch on the Ogbn-arxiv dataset.
Architecture
Message Passing
Layers
Hidden
Dimension
Grid Size Spline Order # Parameters
Training time
(s/epoch)
GIN 2 16 NA NA 5,336 0.035
GIN 2 64 NA NA 27,368 0.041
GIN 4 16 NA NA 6,936 0.037
GIN 4 64 NA NA 52,200 0.056
KAGIN 2 16 1 1 31,232 0.046
KAGIN 2 64 1 1 63,488 0.048
KAGIN 2 16 4 1 54,656 0.059
KAGIN 2 16 4 3 70,272 0.146
KAGIN 4 16 1 1 38,400 0.056
KAGIN 4 64 1 1 116,736 0.072
KAGIN 4 64 4 1 204,288 0.109
KAGIN 4 64 4 3 262,656 0.298
@@ 4.2 Graph Classification
In this set of experiments, we compare the KAGIN model against GIN on standard graph classification benchmark datasets [25]. We experiment with the 7 following datasets: (1) MUTAG, (2) DD, (3) NCI1, (4) PROTEINS, (5) ENZYMES, (6) IMDB-B, (7) IMDB-M. The first $5$ datasets come from bio- and chemo-informatics, while the last $2$ are social interaction datasets.
Experimental setup. We follow the experimental protocol proposed in [8]. Thus, we perform $10$-fold cross-validation to obtain an estimate of the generalization performance of each method, while within each fold a model is selected based on a $90\%/10\%$ split of the training set. We use the splits provided in [8]. We use the Optuna package to select the model that achieves the lowest validation error. We set the number of iterations of Optuna equal to $100$.
For a fair comparison, we set the number of message passing layers of both models to a fixed value for each dataset. Based on preliminary experiments, on MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$. On DD, we set it to $3$. On ENZYMES, we set it to $4$ and finally, on NCI1, we set it to $5$. To produce graph representations, we use the sum operator. The produced graph representations are fed to an MLP (for GIN) or a KAN (for KAGIN) layer which computes the output. We train each model for $1,000$ epochs by minimizing the cross entropy loss. We use the Adam optimizer for model training [19]. We apply batch normalization [17] and dropout [33] to the output of each message passing layer. We also use early stopping with a patience of $20$ epochs. For both models, we choose the number of hidden layers from $\{2,3,4\}$ and the dropout rate from $[0.0,0.5]$. For GIN, we choose the hidden dimension size from $\{8,9,\ldots,256\}$ and the learning rate from $[10^{-5},10^{-2}]$. For KAGIN, we choose the hidden dimension size from $\{2,3,\ldots,128\}$, the grid size from $\{1,2,\ldots,16\}$, the spline order from $\{1,2,\ldots,8\}$ and the learning rate from $[10^{-4},10^{-2}]$. For the IMDB-B and IMDB-M datasets, where there are no node features, we annotate nodes with one-hot encodings of their degrees, up to $35$ (all degrees above $35$ are set equal to $35$). Once the best hyperparemeters are found for a given split, we train $3$ different models on the training set of the split (in order to limit the impact of random initialization) and evaluate them on the test set of the split. This yields $3$ test accuracies, and we compute their average.
@@
Table 3: Average classification accuracy ($\pm$ standard deviation) of the KAGIN and GIN models on the $6$ graph classification datasets.
MUTAG DD NCI1 PROTEINS
GIN 85.09 $\pm$ 5.97 73.68 $\pm$ 3.89 79.57 $\pm$ 1.27 70.26 $\pm$ 3.62
KAGIN
85.45
$\pm$ 7.38
75.46
$\pm$ 4.16
79.72
$\pm$ 1.79
71.78
$\pm$ 3.12
ENZYMES IMDB-B IMDB-M
GIN
53.72
$\pm$ 9.23
73.03 $\pm$ 2.86
49.78
$\pm$ 3.38
KAGIN 42.94 $\pm$ 5.69
73.83
$\pm$ 4.35
49.78
$\pm$ 4.69
@@ Results.
Table 3 illustrates the average classification accuracies and the corresponding standard deviations of the two models on the different datasets. We observe that the two models achieve similar levels of performance. KAGIN outperforms GIN on $5$ out of the $7$ datasets. However, the difference in performance between the two models is very small. This suggests that the two models are similar in terms of expressive power. On ENZYMES, however, KAGIN was found to perform much worse than GIN. Note that ENZYMES consists of more classes ($6$ classes in total) than the rest of the datasets, while ENZYMES and PROTEINS are the only datasets where the nodes of the graphs are annotated with continuous features (on the rest of the datasets, they are annotated with one-hot encodings). We hypothesize that the difference in performance is due to the inability of KANs to handle those continuous features. We thus normalized the node features within each fold by removing the mean of each feature (computed from the training samples) and then dividing by the corresponding standard deviation. We re-conducted the experiment and the average accuracy increased by approximately $6\%$ ($48.77\%$ instead of $42.94\%$). Therefore, it turns out that in some settings it might be harder for KAN layers to handle continuous features than for MLPs.
Training times.
We give in Table 4 an overview of the training times for different GIN and KAGIN architectures. We provide the number of parameters of each architecture. We notice that, for the same number of parameters, KAN is slower than its MLP counterpart. This is particularly sensitive to grid size and spline order. This makes sense since, using splines, each parameter involves more complex computations than the usual multiplication/summation of traditional MLP neurons. Moreover, some of this performance difference might come from how optimized the implementation is. It would be an interesting next step to study the relation of size to performance for KANs and MLPs, since, intuitively, the expressivity of splines should allow for smaller networks.
Table 4: Training time per epoch on NCI1 dataset (i. e., only the forward and backward pass, looped over all batches, averaged over $50$ epochs). All compared models consist of $5$ message passing layers, and of $2$ hidden layers in their respective MLP/KAN. Batch size is set to $128$.
Architecture
Hidden
Dimension
Grid Size Spline Order # Parameters
Training time
(s/epoch)
GIN 16 NA NA 3,522 0.368
GIN 64 NA NA 44,802 0.372
GIN 256 NA NA 670,722 0.433
GIN 512 NA NA 2,652,162 0.742
KAGIN 16 1 1 6,752 0.474
KAGIN 32 1 1 21,696 0.475
KAGIN 64 1 1 76,160 0.475
KAGIN 256 1 1 1,091,072 0.732
KAGIN 16 1 3 10,048 0.658
KAGIN 32 1 3 32,384 0.659
KAGIN 64 1 3 113,920 0.792
KAGIN 256 1 3 1,635,328 2.159
KAGIN 16 4 1 11,696 0.476
KAGIN 32 4 1 37,728 0.478
KAGIN 64 4 1 132,800 0.485
KAGIN 256 4 1 1,907,456 1.074
KAGIN 16 4 3 14,992 0.662
KAGIN 32 4 3 48,416 0.726
KAGIN 64 4 3 170,560 1.005
KAGIN 256 4 3 2,451,712 3.092

### v3

@@ 3.1 The KAGIN Layer
To achieve its maximal expressivity, the GIN model relies on the MLP architecture and its universal approximator property. Since KAN is also a universal function approximator, we could achieve the same expressive power using KANs in lieu of MLPs. We thus propose the KAGIN model, defined as:
@@
With theoretically-sound KANs (i. e., with arbitrarily complex components), this architecture is exactly as expressive as the vanilla GIN model with infinite layer width. In practice, we employ two different families of functions to serve as the base components, leading to two different layers: BS-KAGIN, based on B-splines, as in the original paper; and RBF-KAGIN, based on radial basis functions (RBFs), which were proposed as a more computationally-efficient alternative. While these do not guarantee universal approximation, the empirical results in the original paper demonstrate the great expressive power of KANs [12], motivating our empirical study.
@@ 3.2 The KAGCN Layer
GCN-based architectures have achieved great success in node classification tasks. While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some of the datasets does not facilitate a fair estimation of KANs’ potential in this context. To this end, we also propose a variant of the GCN model. Specifically, we substitute the parameters and ReLU function of the standard GCN [15] model with a single KAN layer (defined in Equation 6) to obtain the KAGCN layer:
@@
where $\Phi^{(\ell)}$ denotes a single KAN layer. In the familiar matrix formulation, where $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with self-loops and $\tilde{\mathbf{D}}$ the diagonal degree matrix of $\tilde{\mathbf{A}}$, the node update rule of KAGCN can be written as $\mathbf{H}^{(\ell)}=\Phi^{(\ell)}\Big{(}\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(\ell-1)}\Big{)}$ where the different rows of $\mathbf{H}^{(\ell)}$ store the representations of the different nodes of the graph. We propose both a BS-KAGCN and a RBF-KAGCN, where $\Phi$ is based respectively on B-splines and RBFs.
@@ 4 Empirical Evaluation
In this section, we compare both KAGIN and KAGCN models against the GIN and GCN models in the following tasks: node classification, graph classification, and graph regression. The code for reproducing the results is available at anonymized. All models are implemented with PyTorch [32]. For KAN layers, we rely on publicly available implementations for B-splines  [^1] https://github.com/Blealtan/efficient-kan and for RBF  [^2] https://github.com/ZiyaoLi/fast-kan. For a given task, all models have consistent architectures, differing as little as possible.
@@ 4.1 Node classification
Datasets
To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer [15] and Ogbn-arxiv [33]) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks. The homophilic networks are already split into training, validation, and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross-validation indices.
@@
We perform two types of experiments. First, we compare our KAGCN model on the homophilic datasets to the GCN proposed in the state of the art approach in [34]. The results are shown in Table 1. One can observe that the classical GCN outperforms the KAN-based model in 2 out of 3 datasets, albeit with a narrow difference and confidence overlaps in all of the comparisons. Note that current setting is built to favor the classical GNNs, as it relies on techniques such as dropout, jumping knowledge, batch normalization and preliminary linear transformations. Hence, although the MLP-based GCN enjoys a plethora of assistive techniques examined in the past decades, the KAN-based counterpart exhibits a close performance. We should note here that these assistive techniques have not been developed yet for KAN-based architectures, and our results aim at motivating the development of such techniques.
Table 1: Average classification accuracy ($\pm$ standard deviation) and model parameters based on Luo et al.[34] using Fastkan, KAGCN, GCN on the Cora, CiteSeer, and Arxiv datasets.
Model Cora CiteSeer Arxiv
Accuracy (%) Params Accuracy (%) Params Accuracy (%) Params
RBF-KAGCN 81.72 $\pm$ 0.79 3,714,217 68.16 $\pm$ 0.61 6,258,808 71.39 $\pm$ 0.28 813,647
BS-KAGCN 80.84 $\pm$ 1.61 3,702,144 70.04 $\pm$ 1.66 4,317,952 70.58 $\pm$ 0.20 1,010,432
GCN 81.60 $\pm$ 0.40 3,262,983 71.60 $\pm$ 0.40 6,221,830 71.74 $\pm$ 0.29 2,591,784
Table 2: Hyperparameter ranges for node classification
Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order
GIN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[8,1024\right]$ NA NA
BS-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[1,8\right]$ $\left[1,4\right]$
RBF-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[2,32\right]$ NA
Our second approach compares the vanilla versions of the neural architectures when we replace the MLP with KAN, as delineated in sections 3. We only use message passing layers, skip connections and an output layer. We do not utilize dropout, batchnorm, jumping knowledge etc. The reason we constrain ourselves to the bare minimum is to isolate and clarify the effect of KAN on the model’s performance. In other words, we aspire to compare the classical MLP-based GNNs as if they were in the same nascent stage as their KAN-based counterparts. Having said that, we perform an auxiliary experiment afterwards, to position the KAGCN in the SOTA and clarify its current utility.
We then evaluate the $10$ models on the test set and report the average accuracy. For the heterophilic datasets, we report the average accuracy across the $10$ folds. In order for all models to exploit the same quantity of information, we fix the number of message-passing layer dataset-wise, with $2$ for Cora and CiteSeer, $3$ for Texas, Cornell, Wisconsin, and Ogbn-arxiv and $4$ for Actor.
For every dataset and model, we tune the values of the hyperparameters using the Optuna package [35]. For each parameterization, we train $5$ models and select the parameterization that yields the lowest mean validation loss (cross-entropy) across all $5$ models. We set the number of iterations of Optuna equal to $100$ trials with a TPE Sampler. We use early stopping with a patience of $50$. Once the best hyperparameter values are found, we evaluate the models on the test set. For the homophilic networks, we initialize and train $10$ different models with the best hyperparameters (using $10$ different random seeds). The ranges for hyperparameters are different for each model. Indeed, for the same number of ”neurons”, a KAN-based layer will have a much larger number of parameters than an MLP-based layer (see Appendix 5.2 for examples). In order to avoid giving an unfair advantage to KAN-based layers, we thus allowed the MLP layers to be much larger, so that the largest possible model in all cases had around the same number of parameters. The hyperparameters ranges are given in table 2.
@@
Model Cora Citeseer Ogbn-arxiv Cornell Texas Wisconsin Actor
GIN 67.48$\pm$6.53 61.04$\pm$6.42 35.68$\pm$12.9 41.70$\pm$1.70 45.14$\pm$6.32 43.24$\pm$2.68 26.43$\pm$0.86
BS-KAGIN 76.28$\pm$0.98 63.42$\pm$2.06 30.53$\pm$5.59 48.92$\pm$1.08 62.30$\pm$0.23 50.78$\pm$1.54 28.00$\pm$0.37
RBF-KAGIN 72.34$\pm$3.74 44.26$\pm$12.0 21.77$\pm$13.2 47.49$\pm$1.85 59.92$\pm$1.88 51.73$\pm$1.51 28.91$\pm$0.40
GCN 75.00$\pm$4.11 62.35$\pm$3.87 40.60$\pm$6.75 41.27$\pm$1.96 55.32$\pm$3.27 52.02$\pm$1.12 28.64$\pm$0.24
BS-KAGCN 78.99$\pm$0.80 66.81$\pm$1.25 52.67$\pm$4.07 48.73$\pm$1.58 58.70$\pm$0.52 56.18$\pm$1.09 27.07$\pm$0.27
RBF-KAGCN 63.59$\pm$7.10 48.00$\pm$8.32 54.56$\pm$3.99 45.86$\pm$2.10 58.43$\pm$1.21 53.35$\pm$1.38 27.10$\pm$0.31
@@ Results.
The results are given in Table 3. We see that KAN-based variants beat MLP-based models on all datasets, while the B-splines-based variant outperforms both other models on $5/7$ datasets. We also observe that RBF-KAGIN seems to exhibit a high instability, especially on Citeseer and Ogbn-arxiv, where its performance is much lower than that of other models, and its variance a lot higher. We notice that the simpler GCN-based models are more efficient than the more complex GIN-based models on all three homophilic datasets, while the opposite is true on heterophilic datasets. This is coherent with the inherent complexity discrepency between both families of datasets [36].
Overall, B-splines are more effective for both graph learning models, the difference being clear mostly in homophilic networks. Overall we contend that KAN has a positive impact on both, the GIN and the GCN architecture. Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e. g., GCN’s low accuracy on heterophilic networks, however, it improves the models in the majority of the datasets.
@@ 4.2 Graph Classification
In this set of experiments, we compare the KAGIN models against GIN on standard graph classification benchmark datasets [37]. We experiment with the following $7$ datasets: (1) MUTAG, (2) DD, (3) NCI1, (4) PROTEINS, (5) ENZYMES, (6) IMDB-B, (7) IMDB-M. The first $5$ datasets come from bio- and chemo-informatics, while the last $2$ are social interaction datasets.
Experimental setup. We follow the experimental protocol proposed in [38]. We kept our architecture simple. First, a number of message-passing layers extract node representation, whose model is either an MLP (resp. a KAN). We then use sum pooling to obtain graph representations. Finally, an MLP (resp. a KAN) with same architecture as those found in the message-passing layer, is used as a classifier. A softmax is used for predicting class probabilities. Thus, we perform $10$-fold cross-validation, while within each fold a model is selected based on a $90\%/10\%$ split of the training set. We use the pre-defined splits provided in [38]. We use the Optuna package to select the model that achieves the lowest validation cross-entropy loss over $100$ iterations. For a fair comparison, we set the number of message-passing layers of all architectures to a fixed value for each dataset. On MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$. On DD, we set it to $3$. On ENZYMES, we set it to $4$ and finally, on NCI1, we set it to $5$. We train each model for $1,000$ epochs (early stopping with a patience of $20$ epochs) by minimizing the cross entropy loss. We use the Adam optimizer for model training [39]. We apply batch normalization [40] and dropout [41] to the output of each message-passing layer. The ranges for hyperparameters search are given in table 4. Once the best hyperparameters are found for a given split, we train $3$ different models on the training set of the split and evaluate them on the test set of the split. This yields $3$ test accuracies for this split, and we compute their average to obtain the performance for this split.
Model Learning Rate Hidden Layers Hidden Dim. Dropout Grid Size Spl. Order
GIN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,512\right]$ $\left[0,0.9\right]$ NA NA
BS-KAGIN/KAGCN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[0,0.9\right]$ $\left[2,32\right]$ $\left[1,4\right]$
RBF-KAGIN/KAGCN $\left[1e-4,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[0,0.9\right]$ $\left[2,16\right]$ NA
Table 4: Hyperparameter ranges for node classification
@@
Table 5: Average classification accuracy ($\pm$ standard deviation) of the KAGIN and GIN models on the $7$ graph classification datasets.
MUTAG DD NCI1 PROTEINS ENZYMES IMDB-B IMDB-M
GIN 85.26 $\pm$ 7.30 74.78 $\pm$ 4.65 78.32 $\pm$ 2.49 71.01 $\pm$ 2.81 53.28 $\pm$ 1.23 73.13 $\pm$ 3.97 50.27 $\pm$ 4.01
BS-KAGIN 85.08 $\pm$ 6.07 75.37 $\pm$ 3.37 79.44 $\pm$ 1.67 74.09 $\pm$ 3.60 42.72 $\pm$ 8.94 73.37 $\pm$ 4.03 49.91 $\pm$ 3.47
RBF-KAGIN 85.48 $\pm$ 5.36 75.66 $\pm$ 3.08 78.33 $\pm$ 1.81 73.29 $\pm$ 3.56 41.17 $\pm$ 8.32 71.97 $\pm$ 3.81 48.76 $\pm$ 4.53
GCN 56.34 $\pm$ 15.1 63.24 $\pm$ 3.82 60.42 $\pm$ 1.01 60.46 $\pm$ 2.06 22.06 $\pm$ 5.58 68.70 $\pm$ 3.63 46.24 $\pm$ 4.83
BS-KAGCN 75.55 $\pm$ 8.14 73.31 $\pm$ 4.45 71.59 $\pm$ 3.31 75.55 $\pm$ 3.62 57.67 $\pm$ 8.03 70.83 $\pm$ 3.95 50.00 $\pm$ 4.41
RBF-KAGCN 73.52 $\pm$ 11.6 68.30 $\pm$ 4.53 63.95 $\pm$ 7.80 74.18 $\pm$ 3.64 59.39 $\pm$ 8.03 72.60 $\pm$ 4.61 48.53 $\pm$ 4.03
@@ Results.
Table 5 illustrates the average classification accuracies and the corresponding standard deviations of the three models on the different datasets. We observe that GIN and both KAGIN  models achieve similar levels of performance on all datasets except ENZYMES. In particular, RBF-KAGIN  and BS-KAGIN  both have the best accuracy on $2$ datasets each, and GIN on $1$ dataset. Regarding the GCN variants, we observe that the KAN-based variants now have a significant advantage over the linear GCN, confirming that a single KAN layer is more expressive than a simple fully-connexted layer. BS-KAGCN  and RBF-KAGCN  obtain the best overall accuracy on $1$ dataset each, while the vanilla GCN performs often significantly poorly.
These experiments show that, on a graph-level task, which requires significantly more power than node-level tasks, KAN-based models seem to perform better than their MLP-based counterparts. While this difference is not significant among GIN-like architectures, it appears that KAN-based GCNs benefit from the higher expressivity in a single layer, and attain performances much closer to their GIN-like counterparts, sometimes even outperforming them.